7.
If y varies directly with x, and y = 1020 when x = 600, find the constant of variation k.




Answer by venugopalramana(3286)

If y varies directly with x, and y = 1020 when x = 600, find the constant of variation k.
Y VARIES DIRECTLY AS X MEANS
Y=K*X...WHERE K IS A CONSTANT
WE GOT.....y = 1020 when x = 600....SO...
1020=K*600
K=1020/600=17